OpenStudy (anonymous):

Help factoring 10-29a+10a^2

OpenStudy (shubhamsrg):

100 = 25*4 29 = 25+4

OpenStudy (anonymous):

how would i write that as an answer

OpenStudy (anonymous):

would have to put in parentheses how would i write that?

OpenStudy (anonymous):

So you know your final answer needs to look like :\\[(5a-**)(2a-**) \] to get your first term back. From there think about the mutiples for the last term which are 10*1 and 5*2 so insert them in to that equation, multliply through, and see what works. So you get \[(5a-10)(2a-1) = 10a^2 -25a +10\] or \[\[(5a-5)(2a-) = 10a^2 -29a +10\]\] So a little guess and check is the last step. Just remember the factor by grouping method

OpenStudy (anonymous):

it said the correct answer is (2a-5)(5a-2) :( im so confused

OpenStudy (anonymous):

I'm sorry. The answer that you just gave is correct. I just didn't check my work too carefully, but what I told you still applies. Use the factor by grouping method and then a little guess and check. Sorry for the confusion. Hope this helps, ask more if you are still confused

OpenStudy (anonymous):

its ok im just confused in general :) it just changed my question up

OpenStudy (anonymous):

Ask and I will do my best to make things clearer